Can a body have a charge of 0.8X10^(-19)C ?

Grade:12

9 Answers

Prajwal kr
10 years ago

No, as the minimum possible is 1.6 x 10^-19, which is the charge of an electron.

rest all
10 years ago

no it cannot have this charge b''coz charge is quantizedwhich means it exist in whole no multiples of 1.6*10^(-19)

anurag hundal
10 years ago

the least charge a body can have(without considering quarks) is -1.6 x 10^-19

Akash Kumar Dutta
10 years ago

not poss since electronic charge=2 times the mentioned charge

a body cant have charge without the integral multiples of the charge of elec.

but quarks can have a charge=q/3...q=charge of electron

Tamana Sehgal
10 years ago

 q=ne

=> n=q/e = 0.8X10^(-19)/ 1.6X10^(-19) = 1/2

As, n is in fraction which is not possible. Hence, a body can''t have a charge of 0.8X10^(-19).

Abhishekh kumar sharma
10 years ago

no

Anangaraj Saikia
5 years ago
No it cant. A charged body cant have a charge of 0.8×10^(-19) bcoz charge can only be quantized in terms of n in which n stands for any integral values i.e, 0,1,2,... and the formula is given by q=ne.
To check whether the given charge is possible or not just divide the given charge by the charge of an electron.
The answer to the given question is 0.5 so it is not possible.
Thank you
 
Rahul
4 years ago
No,because according to quantisation of charge charge cannot take arbitary values but only values that are integral multiples of the fundamental charge
ankit singh
3 years ago
 

no it cannot have this charge b''coz charge is quantizedwhich means it exist in whole no multiples of 1.6*10^(-19)

 
 

the least charge a body can have(without considering quarks) is -1.6 x 10^-19

 
 

not poss since electronic charge=2 times the mentioned charge

a body cant have charge without the integral multiples of the charge of elec.

but quarks can have a charge=q/3...q=charge of electron